derive the expression for the total energy of an electron in the nth orbit


1
Expert's answer
2021-04-23T07:31:31-0400


Knetic energy of electron in nth orbit can be given as -

"KE=\frac{1}{2} mv_n2...........(1)"


Also, we know that,

"vn=\frac{e^2}{2nh}ε°"


∴ By putting the value of Vn in equation (i), we get -

"KE = \frac{1}{2}m(\frac{ e^2}{2nh})^2εο"


"K.E = \frac{me^4}{8n^2h^2} (εο)^2"


In the same way,

Potential energy "(P.E.) = \u221214\u03c0\u03b5^\u2218e \u00d7 2rn"

We know,

"rn = \u03b5^\u2218n^2h^2\u03c0me^2"


Now, Potential energy =

"PE = -\frac{e^2}{4}ε^∘πme^2ε^∘n^2h^2.........(2)"


Hence, "PE = −\frac{me^4}{4}ε^∘2n^2h^2"


Now, Total Energy(En) = Kinetic Energy + Potential Energy

"En = \frac{me^4}{8}ε^∘2n^2h^2–\frac{me^4}{4}ε^∘2n^2h^2"


"∴ En =\frac{me^4}{8}ε^∘2n^2h^2"



Putting the values of m, e, and h in above equation we get,

"m= 9.1 \u00d710^{-31}"


"e = 1.6 \u00d7 10^{-19}"


"\u03b5^o = 8.85 \u00d710^{-12}"


"h = 6.62 \u00d7 10^{-34}"


  we get,

∴ Total Energy of nth orbit = – 13.6/n2 eV